Bigger mammals tend to carry their young longer before birth. The length of pregnancies in horses from conception to birth varies according to a roughly normal distribution with a mean of 336 days and a standard deviation of 3 days.

iii. What is the length of pregnancy below which the shortest 10% of births occur?

with working please
Answer #1

= 336

= 3

Shortest 10% corresponds to area = 0.50 - 0.10 = 0.40 from mid value to Z on LHS.

Table of Area Under Standard Normal Curve gives Z = - 1.28.

So,

we get:

Z = - 1.28 = (X - 336)/3

So,

X = 336 - (3 X 1.28)

= 332.16

So,

Answer is:

332.16